OpenStudy (anonymous):

what is 3 over ten of 20

OpenStudy (anonymous):

3/10 x 20?

OpenStudy (anonymous):

sry but the site wont allow me to show u wat to do so u have to continue it from there and ill correct u :)

OpenStudy (anonymous):

thx

geerky42 (geerky42):

\[\dfrac{3}{10}\cdot 20 = \dfrac{3·20}{10} = \dfrac{60}{10}\]

OpenStudy (anonymous):

@geerky42 60/10 can be simplified as 6